OpenStudy (anonymous):

Am I correct? A circle has its center at (1, 4) and a radius of 2 units. What is the equation of the circle? (x - 1)2 + (y - 4)2 = 4 (x + 2)2 + (y + 4)2 = 2 (x + 1)2 + (y - 4)2 = 4 (x - 1)2 + (y - 4)2 = 2 * I think the answer is the 4th choice*

OpenStudy (anonymous):

The circle formula is (k-x)^2 + (k -y)^2 = r^2 For your formula, the answer is A

OpenStudy (anonymous):

ok thanks yeah that radius part is what i was not sure of

OpenStudy (zepp):

^

OpenStudy (zepp):

The right side is r^2, they just expand it, 2^2=4

OpenStudy (anonymous):

thanks :)